# Towards Quantifying Long-Range Interactions in Graph Machine Learning: a Large Graph Dataset and a Measurement

### v1

@@ Abstract
Long-range dependencies are critical for effective graph representation learning, yet most existing datasets focus on small graphs tailored to inductive tasks, offering limited insight into long-range interactions. Current evaluations primarily compare models employing global attention (e.g., graph transformers) with those using local neighborhood aggregation (e.g., message-passing neural networks) without a direct measurement of long-range dependency. In this work, we introduce City-Networks, a novel large-scale transductive learning dataset derived from real-world city roads. This dataset features graphs with over $10^{5}$ nodes and significantly larger diameters than those in existing benchmarks, naturally embodying long-range information. We annotate the graphs using an eccentricity-based approach, ensuring that the classification task inherently requires information from distant nodes. Furthermore, we propose a model-agnostic measurement based on the Jacobians of neighbors from distant hops, offering a principled quantification of long-range dependencies. Finally, we provide theoretical justifications for both our dataset design and the proposed measurement—particularly by focusing on over-smoothing and influence score dilution—which establishes a robust foundation for further exploration of long-range interactions in graph neural networks.
@@ 1 Introduction
Graphs are a widely used mathematical abstraction across nearly every branch of science. They are particularly effective for modeling the intricate and non-uniform interactions found in real-world data, where nodes stand in for objects and edges depict their connections. The growing recognition of the versatility of graph representations has sparked intense interest in Graph Neural Networks (GNNs) (Scarselli et al., 2009), driving innovation in deep learning for both geometric and graph-centric applications. Most GNNs exchange information between one-hop neighbors per layer to build node representations, hindering their ability to model long-range interactions. Indeed, the majority of GNN architectures have remained shallow compared to their counterparts in the Large Language Model (LLM) and Computer Vision (CV) literature, arguably due to the small world effect (Watts & Strogatz, 1998) observed in most citation and social networks (Yang et al., 2016b), which have traditionally been the focus of graph representation learning research. Generally speaking, despite comprising thousands of nodes, such networks have short average path lengths and high clustering coefficients, making it possible to process a high proportion of the network with only a few message-passing layers.
@@
To better understand long-range interactions, recent efforts, such as the Long Range Graph Benchmark (LRGB(Dwivedi et al., 2023), have introduced alternative graph datasets. However, two significant challenges remain. First, the LRGB datasets focus on inductive learning tasks with relatively small graphs of around $500$ nodes, while currently there is no benchmark for long-range dependencies on large graphs for transductive learning in the literature. This is a critical gap since applying standard Graph Transformers, which are expected to better capture long-range interactions, to large graphs is significantly more challenging compared to small-graph inductive learning tasks due to the computational complexity of global attention (Borde, 2024). Second, to justify the long-range dependencies in their datasets, the authors of LRGB primarily focused on comparing classical GNNs with Graph Transformers, where they empirically observed performance gaps which are then associated with the capability of long-range signal handling. However, conclusions that are solely derived from comparisons of downstream performance gaps may not be reliable as they can be largely influenced by hyperparameter tuning as shown in Tönshoff et al. (2023), leading to an ambiguous assessment of the long-range interactions.
To address these limitations, we propose City-Networks, a transductive learning dataset based on large-scale city road networks with long-range annotations, and introduce a measurement to quantify long-range dependency. We then provide theoretical justification of both. Our contributions can be summarized as follows:
@@
  • Our dataset consists of four city road networks with a topology distinct from those commonly found in the literature. In particular, it features grid-like large graphs with up to $500k$ nodes and diameters of up to $400$, where the node labels are annotated based on an approximation of eccentricity, which inherently requires long-range dependency for its calculation. To the best of our knowledge, this is the first large-graph dataset designed for testing long-range dependencies in graph representation learning.
@@
  • We design a methodology to empirically test classical GNNs and Graph Transformers on our dataset by varying the number of model layers and sampled neighborhood hops. The results show that communication with neighbors from distant hops consistently improves the performance of all models, supporting the presence of long-range signals in our dataset.
@@
  • To quantify such long-range dependency, we further introduce a model-agnostic measurement that considers the per-hop influence of a given node’s neighbors on its prediction, which is estimated by the aggregated $\ell_{1}$-norm of the Jacobian from a trained model at each hop around the given node. Interestingly but not surprisingly, we observe that distant hops exert a greater influence on our City-Networks compared to other social networks in the transductive learning literature.
@@
  • We theoretically justify the graph structure in our dataset from a spectral perspective on over-smoothing, whose rate we link to the algebraic connectivity and diameter of the graph. Similarly, we relate our proposed measurement to the concept of influence as defined in the literature and study the dilution of mean influence score in grid-like graphs.
2 Datasets: City Road Networks
As shown in Figure 1, our proposed City-Networks is based on street maps of four cities [^1] Code for reproducing our dataset and experiments can be found at github.com/LeonResearch/City-Networks. —Paris, Shanghai, Los Angeles, and London, which are obtained by querying OpenStreetMap (Haklay & Weber, 2008) with OSMnx (Boeing, 2024). In the following sections, we first compare the characteristics of our proposed networks with existing benchmarks, and then discuss the construction of node features and labels in detail, after which we provide a task description with intuitive visualizations and explanations. Readers are also referred to Appendix A for more details regarding our dataset.
Network statistics.
We consider a city network inclusive of all types of roads in the city (e.g., drive, bike, walk, etc.), where nodes in the graph represent road junctions, and edges represent the road segments. After obtaining the raw data, we post-process the city networks by 1) making the graphs undirected so that they only represent connectivity between junctions (see Appendix A for details); 2) retaining the largest connected component in each city and summarize their network statistics in Table 1. Compared to citation networks from Planetoid (Yang et al., 2016a) and OGB (Hu et al., 2020), our city networks exhibit much lower maximum and average degrees with low standard deviations. At the same time, they have much larger diameters with minimum estimation ranging from 100 to 400, whereas the citation networks typically have diameters of around 20 to 30. Compared to super-pixel graphs from LRGB [^2] In the case of LRGB each dataset has multiple graphs; we compute average statistics for 1,000 samples., our city networks have a much larger graph size of more than $10^{5}$ nodes with large diameters, and meanwhile maintain a relatively low clustering coefficient and transitivity. We refer readers to Appendix A.1 for a more detailed explanation of these statistics and our estimation methods.
Table 1: Summary statistics of our City-Networks compared to LRGB, Planetoid, and OGB.
Dataset City-Networks (ours) LRGB Planetoid OGB
Statistics Paris Shanghai L.A. London PascalVOC COCO Cora CiteSeer ogbn-arxiv
# Nodes $114k$ $184k$ $241k$ $569k$ $479$ $477$ $2.7k$ $3.3k$ $169k$
# Edges $183k$ $263k$ $343k$ $759k$ $1.3k$ $1.3k$ $5.3k$ $4.6k$ $1.16m$
Avg. degree $3.2$ $2.9$ $2.8$ $2.7$ $5.7$ $5.6$ $3.9$ $2.7$ $13.7$
Std. degree $0.8$ $1.0$ $1.0$ $1.0$ $1.2$ $1.2$ $5.2$ $3.4$ $68.6$
Max. degree $15$ $8$ $9$ $10$ $10$ $10$ $168$ $99$ $13k$
Avg. cluster coef. $0.03$ $0.04$ $0.04$ $0.04$ $0.43$ $0.43$ $0.24$ $0.14$ $0.23$
Transitivity $0.03$ $0.04$ $0.05$ $0.05$ $0.40$ $0.40$ $0.09$ $0.13$ $0.02$
Diameter est. $121$ $123$ $171$ $404$ $28$ $27$ $19$ $28$ $25$
Node homophily $0.70$ $0.75$ $0.75$ $0.76$ $0.92$ $0.90$ $0.81$ $0.73$ $0.65$
Real-world node features.
The node features in our dataset are derived from real-world features provided by OpenStreetMap for both road junctions and road segments. Specifically, the original node features consist of longitude and latitude of the road junction, the adjacent road count, as well as type of land use at the coordinate (e.g., residential, industrial, forest, etc.). For edge features, real-valued features like road length and speed limit are available, as well as binary features like one-way and reversed and categorical features such as road types (e.g., primary, service, residential, etc.). After one-hot encoding binary and categorical features, we obtain a total of $12$ node features and $25$ edge features. Next, we apply a simple neighborhood aggregation that transfers edge features to node features by averaging the features of incidental edges and then concatenating them to the features of the focal node. These new node features represent, for instance, the average speed limit around a road junction or the probability of finding a residential road nearby. As a result, the final dataset contains $37$ node features after processing, where the original edge attributes are then discarded.
Node labels based on eccentricity.
To create class labels that require long-range dependency, we compute an approximation of eccentricity $\varepsilon(v)$ in network science (Newman, 2018), which measures the maximum distance from a node $v$ to all other nodes in the graph $\mathcal{G}=(V,E)$ by the following expressions:
$\varepsilon(v)=\max_{u\in V}~{}\rho_{w}(v,u),\quad\rho_{w}(v,u)=\min_{\pi\in P(v,u)}~{}\sum_{e\in\pi}w(e),$ (1)
where $\rho_{w}(v,u)$ is the weighted shortest path distance from node $v$ to node $u$, $P(v,u)$ is the set of all possible paths from node $v$ to node $u$, and $w(e)$ denotes the edge weight (we use road length here) for edge $e$. However, obtaining $\varepsilon(v)$ for all nodes $v\in V$ requires computing all pair-wise shortest paths in the network, which has a computational complexity of $\mathcal{O}(|V|^{3})$ using the Floyd-Warshall algorithm (Floyd, 1962) hence prohibitively expensive for our large-scale networks. Instead of focusing on the entire graph, for each node $v$, we only consider its $16$-hop neighborhood, that is, an ego-network centered at $v$, and compute an approximation, $\hat{\varepsilon}(v)$, based on this subgraph. Although this might seem like a computational limitation in principle, it actually allows us to know a priori that the long-range signal should be highly correlated with the 16th hop, as we later observe in Section 3. To design a classification task, we split the eccentricities for all nodes into $10$ quantiles which we use as node labels. Note that, in contrast, LRGB uses the semantic context of an image after segmentation as labels and treats the synthetic super-pixels as graph nodes, whereas City-Networks is derived from real-world road networks and is therefore a more realistic and practical task that studies the accessibility of different neighbourhoods in a city. We refer readers to Appendix A.2 for more information on our feature construction and labeling approach.
Visualizing labels and long-range dependency in City-Networks.
As shown in Equation (1), calculating node eccentricity $\varepsilon(v)$ (or its approximation $\hat{\varepsilon}(v)$) naturally requires information from distant-hop neighbors, and it carries a practical meaning which relates to the general accessibility of node $v$ in the network. For instance, for locations in the city center, where the network is more interconnected, eccentricities tend to be lower than in rural areas with sparser road connections. To intuitively understand the node labels, we provide two visualizations of Huangpu, Shanghai, and Pasadena, Los Angeles in Figure 2. Interestingly, since we are using the approximation $\hat{\varepsilon}(v)$ based on node $v$’s $16$-hop neighborhood, nodes on major transportation routes such as freeways and highways will tend to have larger eccentricities than those in populated areas. This is because reaching a node’s 16th-hop neighbors from a highway junction often requires traveling a much longer distance compared to road junctions in a downtown area. Based on Figure 2, it is also clear that a significant part of the graph topology is grid-like and possibly quasi-isometric to a lattice, as we will later discuss in Section 6. Note that the graph homophily scores, as shown in Table 1, will not be necessarily low in this case, since nodes with similar $\hat{\varepsilon}$ tend to cluster together.
@@ 2 Datasets: City Road Networks
Figure 2: Visualizations of node eccentricity estimations on (a) Huangpu, Shanghai and (b) Pasadena, L.A., where darker colors indicate smaller eccentricity values, i.e., nodes that are easier to access.
@@
Large networks present a significant challenge for both GNNs and Graph Transformers since they must process features from a node’s distant neighbors—such as their coordinates, land use, and surrounding road characteristics—to accurately determine the target node’s eccentricity. In particular, this requirement creates a fundamental trade-off for GNNs: while increasing the number of layers helps propagate information from distant hops, it also leads to issues such as over-smoothing (Li et al., 2018; Nt & Maehara, 2019; Rusch et al., 2023), over-squashing (Alon & Yahav, 2021; Topping et al., 2022; Di Giovanni et al., 2023), and potential over-parameterization. Similarly, Graph Transformers, which rely on attention mechanisms with quadratic computational complexity, face scalability challenges when applied to our large-scale city networks with long diameters, making it more difficult to efficiently capture global structural information compared to learning on smaller graphs like those in LRGB.

### v2

@@ Abstract
Long-range dependencies are critical for effective graph representation learning, yet most existing datasets focus on small graphs tailored to inductive tasks, offering limited insight into long-range interactions. Current evaluations primarily compare models employing global attention (e.g., graph transformers) with those using local neighborhood aggregation (e.g., message-passing neural networks) without a direct measurement of long-range dependency. In this work, we introduce City-Networks, a novel large-scale transductive learning dataset derived from real-world city road networks. This dataset features graphs with over $10^{5}$ nodes and significantly larger diameters than those in existing benchmarks, naturally embodying long-range information. We annotate the graphs based on local node eccentricities, ensuring that the classification task inherently requires information from distant nodes. Furthermore, we propose a model-agnostic measurement based on the Jacobians of neighbors from distant hops, offering a principled quantification of long-range dependencies. Finally, we provide theoretical justifications for both our dataset design and the proposed measurement—particularly by focusing on over-smoothing and influence score dilution—which establishes a robust foundation for further exploration of long-range interactions in graph neural networks. Our dataset and measurement are now both available on PyTorch Geometric, and the code for reproducing the experimental results can be found at github.com/LeonResearch/City-Networks.
@@ 1 Introduction
Graphs are a widely used mathematical abstraction across nearly every branch of science. They are particularly effective for modeling the intricate and non-uniform interactions found in real-world data, where nodes stand in for objects and edges depict their connections. The growing recognition of the versatility of graph representations has sparked intense interest in Graph Neural Networks (GNNs) (Scarselli et al., 2009; Wu et al., 2021), driving innovation in deep learning for both geometric and graph-centric applications. Most GNNs, in particular variants based on a message-passing mechanism (Gilmer et al., 2017), exchange information between one-hop neighbors per layer to build node representations. While they found wide success in analysing citation and social networks (Yang et al., 2016b), one significant limitation concerns their capability of handling long-range dependencies, where interactions between distant nodes might be required to solve a task. Most existing datasets are not sufficient in assessing this: for instance, social networks, despite comprising thousands of nodes, often exhibit the small world effect (Watts & Strogatz, 1998) with short average path lengths and high clustering coefficients. The node labels on these graphs usually possess a homophilic pattern where nodes tend to connect with “similar” or “alike” others (McPherson et al., 2001), making it possible to propagate sufficient information for modelling with only a few message-passing layers. On the other hand, while connected nodes tend to have different properties on heterophilic graphs (Zhu et al., 2020; Ma et al., 2022), solving the tasks in those cases do not necessarily require the handling of long-range dependencies (Arnaiz-Rodriguez & Errica, 2025).
@@
Recently, Long Range Graph Benchmark (LRGB(Dwivedi et al., 2023) introduces alternative graph datasets based on super-pixels and molecules. To justify the existence of long-rangeness, the authors compare classical GNNs with Graph Transformers (GTs), which leverage global attentions over the entire graph, and associate the observed performance gaps with the presence of long-range dependencies. However, conclusions that are solely derived from empirical comparisons may not be reliable as they can be largely influenced by hyperparameter tuning (Tönshoff et al., 2024), leading to an ambiguous assessment of the long-range interactions. Moreover, LRGB and other synthetic long-range benchmarks (Bodnar et al., 2021; Zhou et al., 2025) all focus on inductive learning tasks with relatively small graphs—typically on the order of $10$ to $10^{2}$ nodes, while currently there is no long-range benchmark that considers large graphs with real-world topology for transductive learning. This is a critical gap in the literature since applying GTs (Rampášek et al., 2022), which are expected to better capture long-range interactions, to large graphs is significantly more challenging compared to small-graph inductive tasks due to the computational complexity of global attention (Borde, 2024).
We aim to address these limitations in this work, and our main contributions are as follows:
@@
  • We propose City-Networks, a transductive learning dataset that consists of four large-scale city road networks with a topology distinct from those commonly found in the literature. In particular, it features grid-like large graphs with up to $500k$ nodes and diameters of up to $400$, where the labels are annotated based on node accessibility that inherently requires long-range dependency in its calculation. To the best of our knowledge, this is the first large-graph dataset designed for testing long-range dependencies in graph representation learning.
@@
  • We empirically test classical GNNs and GTs on our dataset under different model depths, and compare their behaviors to those on other common graph datasets that are short-ranged, large-scale, heterophilic, or long-range dependent. The results on our datasets, unlike those on the existing datasets, suggest that communication with neighbors from distant hops consistently improves the performance of all models, supporting the presence of long-range signals.
@@
  • To quantify such long-range dependency, we further introduce a model-agnostic measurement that quantifies per-hop influence of a focal node’s neighbors on its prediction, which is estimated by the aggregated $\ell_{1}$-norm of the Jacobian from a trained model at each hop around the focal node. As a validation, we observe that distant hops exert a greater influence on all baseline models in our City-Networks compared to those on other commonly used graph datasets in the literature.
@@
  • Finally, we theoretically justify the graph structure in our dataset from a spectral perspective on over-smoothing, whose rate we link to the algebraic connectivity and diameter of the graph. In addition, we relate our proposed measurement to the concept of influence as defined in the literature, and study the dilution of mean influence score in grid-like graphs to justify our design.
2 The City-Networks Dataset
In this section, we begin by identifying the limitations and challenges in current literature. Then, we characterize how the topology of City-Networks differs from the existing datasets, and proceed to justify the rationale behind our long-range labeling strategy. Lastly, we explain how our dataset addresses the current limitation and discuss the new challenges brought to the field.
Challenges in testing long-range dependency. To design and fairly evaluate graph datasets of long-range dependency, we need to address three research challenges:
  1. 1.
    How can we generate long-range signals in a principled and controllable manner, so that models are required to communicate with sufficiently distant neighbors of a node to predict its label?
  2. 2.
    Beyond the small graphs used in LRGB or other synthetic benchmarks, how can we design long-range signals on large real-world networks to test the scalability of a model?
  3. 3.
    How can we define a principled measure to quantify and compare the level of long-rangeness across different datasets?
We will discuss how our proposed dataset handles the first two challenges in this section, and address the third challenge in Section 4.
Table 1: Statistics of City-Networks compared to common graph datasets, where $d$, $C$, $T$, diamdiam, homohomo represent degree, clustering coefficient, transitivity, diameter, and homophily, respectively.
Dataset #Nodes #Edges avg($d$) std($d$) max($d$) $C$ $T$ diamdiam homohomo
Paris $114k$ $183k$ $3.2$ $0.8$ $15$ $0.03$ $0.03$ $121$ $0.70$
Shanghai $184k$ $263k$ $2.9$ $1.0$ 88 $0.04$ $0.04$ $123$ $0.75$
Los Angeles $241k$ $343k$ $2.8$ $1.0$ 99 $0.04$ $0.05$ $171$ $0.75$
London $569k$ $759k$ $2.7$ $1.0$ $10$ $0.04$ $0.05$ $404$ $0.76$
Cora $2.7k$ $5.3k$ $3.9$ $5.2$ $168$ $0.24$ $0.09$ 1919 $0.81$
ogbn-arxiv $169k$ $1.16m$ $13.7$ $68.6$ $13k$ $0.23$ $0.02$ $25$ $0.65$
Amazon-Ratings $24k$ $93k$ $7.6$ $6.0$ $132$ $0.58$ $0.31$ $46$ $0.38$
PascalVOC-SP $479$ $1.3k$ $5.7$ $1.2$ $10$ $0.43$ $0.40$ $28$ $0.92$
2.1 Large-scale Road Networks with Long Diameters
Real-world network topologies and features. As shown in Figure 1, our City-Networks consists of street maps in four cities: Paris, Shanghai, Los Angeles, and London, which are obtained by querying OpenStreetMap (Haklay & Weber, 2008) with OSMnx (Boeing, 2024). In particular, we consider a city network inclusive of all types of roads in the city (e.g., drive, bike, walk, etc.), where nodes represent road junctions with features like longitude, latitude, land use, etc.; and edges represent road segments with features like road length, speed limit, road types, etc. Next, to facilitate a typical node classification task, we apply a simple neighborhood aggregation that transforms edge features into node features by averaging the features of incidental edges and then concatenating them to the features of the focal node. These new node features represent, for instance, the average speed limit around a road junction or the probability of finding a residential road nearby. As a result, the final dataset contains $37$ node features, where we also make the graphs undirected so that they only represent connectivity between junctions, and retain the largest connected component in each city. We refer readers to Appendix B.1 for a more detailed discussion on our dataset processing approach.
Large graphs with long diameters and low maximum degrees. We can observe from Table 1 that our datasets feature large diameters from $100$ to $400$, which are much longer than those of the common graphs used in the literature. In particular, compared to the super-pixel graphs in PascalVOC-SP with typically $500$ nodes, our datasets have much larger sizes ranging from $100k$ to $500k$. Meanwhile, compared to social networks such as Cora, ogbn-arxiv, and Amazon-Ratings, our City-Networks have smaller average clustering coefficients and much lower maximum degrees. We argue that this distinct graph topology enables us to effectively design learnable long-range signals on graphs, as explained in the next paragraph and further justified theoretically in Section 5.
2.2 Long-range labels based on urban accessibility
Based on the city road networks, we design a real-world task that requires the handling of long-range dependencies. Specifically, we aim to predict how accessible a road junction is based on its own location as well as neighboring characteristics of the urban landscape. This is a useful task especially when it comes to evaluating urban design principles such as the 15-minute city (Abbiasov et al., 2024). To quantify accessibility, we measure the distance one needs to travel from one road junction to its $k$-hop neighbors along road segments in the road network. By design, solving this task requires the model to be able to access information within the $k$-hop neighborhood of each focal node and, by adjusting $k$, we can design a long-range task as desired. Such a notion of accessibility is directly related to node eccentricity $\varepsilon(v)$ in network science (Newman, 2018), which looks at the maximum distance from $v$ to all other nodes in the graph $\mathcal{G}=(V,E)$.
However, using the exact eccentricity as signals is not ideal for two reasons. First, we cannot control the level of long-rangeness, as eccentricity is determined by the entire graph structure and not adjustable by design. In particular, although predicting $\varepsilon(v)$ appears to require distant information based on its definition, in many cases (Figure 6, Appendix B.2), the underlying signal can be well inferred by spatial features (e.g., geographical coordinates) alone. This is because nodes in the city center generally have lower eccentricity due to the grid-like topology in road networks. Second, the calculation of eccentricity requires computing all pair-wise shortest paths in the network, which is typically at cost $\mathcal{O}(|V|^{3})$ and is hence prohibitively expensive for labeling our large-scale networks.
Node classification task based on local eccentricity. With the above consideration in mind, we propose a local eccentricity measure $\hat{\varepsilon}_{k}(v)$ that only considers $k$-hop neighbors $\mathcal{N}_{k}(v)$ of node $v$ when calculating the longest shortest paths with the following expressions:
$\hat{\varepsilon}_{k}(v)=\max_{u\in\mathcal{N}_{k}(v)}~\rho_{w}(v,u),\quad\rho_{w}(v,u)=\min_{\pi\in P(v,u)}~\sum_{e\in\pi}w(e),\vskip-2.84544pt$ (1)
where $\rho_{w}(v,u)$ is the weighted shortest path distance from node $v$ to node $u$, $P(v,u)$ is the set of all possible paths from node $v$ to node $u$, and $w(e)$ denotes the edge weight for edge $e$ (we use road length for this particular purpose, while in all other cases, the graphs are considered unweighted). To design a classification task, we then split $\hat{\varepsilon}_{k}(v)$ for all nodes into $10$ quantiles which we use as node labels for the classification task.
@@ 2.2 Long-range labels based on urban accessibility
Figure 2: Visualizations of node accessibility estimations based on local eccentricity in two sub-regions, where darker colors indicate smaller eccentricity values, i.e., nodes that are easier to access.
We emphasize that this local approach is introduced not only for computational efficiency, but it also brings in two important considerations. First, it creates long-range signals that extend to $k$-hop neighbors by design, i.e., a known “ground-truth” range, hence is directly controllable and testable. Second, it ensures both graph topology and node features contribute to the modelling, avoiding the problem of high correlation between node signal with either the graph structure or node features alone. Indeed, since $k$ and the edge weights used to compute $\hat{\varepsilon}_{k}(v)$ are unknown to the model, it must integrate structural information from distant neighborhoods with their spatial information (geographic location, road type, land use, etc.) to infer the long-range signal, which makes the task non-trivial.
Choice of the long-range level $k$. To assess long-range dependency, $k$ should be sufficiently large to distinguish our setting from short-range tasks (e.g., social graphs with typical message passing around $4$ hops). Thanks to the large grid-like topology with long diameters, we can create long enough node signals based on different local networks. At the same time, $k$ should also not be too large such that (i) the selected neighborhoods maintain local characteristics; and (ii) baseline models under a $k$-layer architecture, which captures the required information, can fit into common GPUs for fair benchmarking purposes. After experimenting with values from $8$ to $32$ (Figure 8, Appendix B.4), we design the task at kk=$16$, which strikes a balance between sufficient long-rangeness and computational efficiency. We refer readers to Appendix B.4 for a more detailed justification of our methods.
Visualizations and interpretations. As shown in Equation (1), calculating $\hat{\varepsilon}_{k}(v)$ naturally requires information from distant neighbors up to $k$ hops, and it carries a practical meaning which relates to the accessibility of node $v$ in the network. As an illustration, we visualize two sub-regions from the city maps in Figure 2 with kk=$16$. We can observe that nodes on major transportation routes such as freeways and highways tend to have larger eccentricities than those in populated areas. This is because reaching a node’s 16-hop neighbors from a highway junction often requires traveling a much longer distance compared to road junctions in a downtown area. It is also clear from Figure 2 that a significant part of the graph topology is grid-like and possibly quasi-isometric to a lattice, as we will later discuss in Section 6. Note that the homophily scores, as reported in Table 1, are reasonably high since nodes with similar $\hat{\varepsilon}_{k}$ tend to cluster together.
@@
New challenges for learning long-range dependencies. The proposed dataset presents a significant challenge for both GNNs and GTs. In general, such long-range dependencies create a fundamental trade-off for graph learning: while increasing the number of layers in GNNs helps propagate information from distant hops, it also leads to issues such as over-smoothing (Li et al., 2018; Nt & Maehara, 2019; Rusch et al., 2023), over-squashing (Alon & Yahav, 2021; Topping et al., 2022; Di Giovanni et al., 2023), and vanishing gradients (Arroyo et al., 2025). On the other hand, GTs, which rely on attention mechanisms with quadratic computational complexity, face scalability challenges when applied to our large-scale city networks compared to learning on smaller long-ranged graphs like those in LRGB (Dwivedi et al., 2023) and other synthetic benchmarks (Zhou et al., 2025).